OpenStudy (anonymous):

How do you write 9,320 in scientific notation

jimthompson5910 (jim_thompson5910):

place a decimal between the first two digits (9 and 3) to get 9.320 you can drop that last 0 because 9.32 = 9.320

jimthompson5910 (jim_thompson5910):

so we have 9.32

jimthompson5910 (jim_thompson5910):

how many spots to the right must we move the decimal point to go from 9.32 to get back to 9,320 ?

OpenStudy (triciaal):

the general formula for scientific notation is a * 10^n where a is >0 and < 10 and n can be positive or negative whole number.

OpenStudy (anonymous):

so would it be 9.32x10 to the 2??

OpenStudy (triciaal):

as @jim_thompson5910 has shown you need to move the decimal point to get the a value you are not changing the number only expressing it differently so if you move the decimal to the right you are multiplying so to keep the balance you divide by 10 t

OpenStudy (triciaal):

you moved 3 places to the left same as divide by 10^3 so to keep the balance its 9.320 * 10^3

OpenStudy (triciaal):

remember when you multiply numbers with the same base add the powers 10^-3 * 10^3 = 10^(-3 + 3 )= 10^0 any number raised to 0 = 1

OpenStudy (triciaal):

it is always better to count off the decimal places moved before dropping the insignificant zero. this was a perfect example of why